# Chiral-split magnons in the S = 1 Shastry-Sutherland model

### v1

@@ I Introduction
In this work, we demonstrate that the collinear Néel phase of the $S=1$ Shastry-Sutherland model possesses altermagnetic character. We establish this both from symmetry considerations and from a direct analysis of magnon dispersions. Our study reveals a unique form of chirality-splitting in the magnon bands: the splitting reverses sign along two different directions in momentum space. Furthermore, we compute the resulting spin Seebeck and spin coefficients using the Kubo formalism and show that these effects remain robust without the need for spin-orbit coupling or applied magnetic fields. Our findings enrich the growing landscape of altermagnetic phenomena and suggest new pathways for controlling magnonic transport in antiferromagnetic systems.
@@ II Model and Methods
where $\mathbf{S}_{i}$ denotes the spin operator on the $i$th site. As illustrated in Fig. 1a $J_{ij}=J$, $J^{\prime}$, and $J^{\prime\prime}$ represent the intra-dimer, the inter-dimer and the inter-chain couplings, which are all assumed to be antiferromagnetic. The single-ion anisotropy is denoted as $D<0$.
@@
The Néel (I) phase (Fig. 1b) of the orthogonal-dimer structure belongs to the plane group $p4g$; the 4-fold axes are at the center and corner of the unit cell, and the 2-fold axes (inversion centers) are at the center of the edges of the cell. We verify the Néel (I) phase to be altermagnetic by checking the identification rules [19]:
@@
where the Nambu spinor $\Psi_{\mathbf{k}}=\left({b^{1}_{\mathbf{k}}},~{}{b^{2}_{\mathbf{k}}},~{}{b^{3}_{\mathbf{k}}},~{}{b^{4}_{\mathbf{k}}};~{}{b_{-\mathbf{k}}^{1}}^{\dagger},~{}{b_{-\mathbf{k}}^{2}}^{\dagger},~{}{b_{-\mathbf{k}}^{3}}^{\dagger},~{}{b_{-\mathbf{k}}^{4}}^{\dagger}\right)^{T}$, with ${b^{l}_{\mathbf{k}}}^{\dagger}$ $(l=1,~{}2,~{}3,~{}4)$ being the annihilation operator defined at the four sites in the unit cell. The Hamiltonian matrix in the reciprocal space takes the form of
@@
where $A(\mathbf{k})$ and $B(\mathbf{k})$ are $4\times 4$ Hermitian matrices:
@@
Figure 2: Magnon band dispersion of the generalized Shastry-Sutherland model near $\mathbf{k}=0$. Red and blue indicate the left and right-handed magnon modes, respectively. The values $J^{\prime}/J=1.2$, $J^{\prime\prime}/J=1$, and $D/J=-0.6$ are used for the parameters.
A distinct chirality, corresponding to the direction of precession of the spin around the axis of the collinear magnetic order, can be assigned to each of the magnon bands. Since the collinear order in the $z$-direction preserves the $z$-component of the total spin $S_{z}=\sum_{i}S_{i}^{z}$ with $i$ running over all sites, we use $\vartheta=\langle S^{z}\rangle/\hbar$ to specify the chirality of the bands [23]. In particular, we have $\vartheta=-1$ for the right-handed and $\vartheta=+1$ for left-handed magnon modes. In other words, magnon modes with chirality $\vartheta=\pm 1$ carry ±$\pm 1$ spin angular momentum along the $z$-direction. Since there are distinct upper and lower bands in the dispersion with opposite chiralities in each, we label the bands as $E^{u}_{\pm 1}(\mathbf{k})$ and $E^{l}_{\pm 1}(\mathbf{k})$, respectively, where the subscript denotes the chirality $\nu$ of the band (Fig. 2).
@@
where $\theta$ is the angle between the temperature gradient $\nabla T$ and the $\hat{\mathbf{x}}$. Based on the Kubo formula [24, 25, 26, 27, 28, 29], we can calculate the magnon conductivity tensor $\sigma_{mn}=\sum_{\mu}\sigma_{mn}^{\mu}$ where $\sigma_{mn}^{\mu}$ is the contribution from the mode $\mu\in\{E^{u}_{\pm 1},~{}E^{l}_{\pm 1}\}$
@@
where $\bm{\sigma}$ is the original thermal conductivity tensor, and $\bm{\sigma}^{\prime}$ gives the components in the rotated frame. In this rotated frame, the diagonal component $\sigma^{\prime}_{x^{\prime}x^{\prime}}$ corresponds to the spin Seebeck response (along the gradient) [30], and the off-diagonal component $\sigma^{\prime}_{y^{\prime}x^{\prime}}$ corresponds to the spin Nernst response [31, 32, 33](transverse to the gradient).
@@ III Results
Figure 3: (a) Energy difference between upper bands with right and left-handed chirality, (b) energy difference between lower bands with right and left-handed chirality, (c) chirality-split dispersion relation of magnon along the path $\Gamma\to M\to X\to\Gamma\to Y\to X$ shown in the inset. We first trace the path $\Gamma\to M\to X\to\Gamma$ (blue arrows), and then $\Gamma\to M\to X$ (red arrows). Red and blue denote the left and right-handed magnon modes, respectively. Here we used $J^{\prime}/J=1.2$, $J^{\prime\prime}/J=1$, and $D/J=-0.6$. (d) Variation of spin Seebeck ($\sigma_{\parallel}$) and spin Nernst ($\sigma_{\perp}$) coefficients for different values of $\theta$. Here, in addition, we used $J/k_{B}=100\text{ K}$ and $T=25\text{ K}$.
The magnon dispersion is shown in Fig. 2. We have used the parameter values $J^{\prime}/J=1.2$, $J^{\prime\prime}/J=1$, and $D/J=-0.6$ unless otherwise stated. There are four bands: two with opposite chiralities with higher energy, $E^{u}_{\pm 1}$, and two with opposite chiralities with lower energy, $E^{l}_{\pm 1}$, with a strict energy gap between these two sectors. We will refer to these two pairs of bands as the lower and the upper bands, respectively.
We observe a $C_{2}$ symmetry in all of the magnon bands, as expected. To observe the chirality split, we calculated the energy difference between the upper bands $\Delta^{u}(\mathbf{k})=E^{u}_{-1}-E^{u}_{+1}$ and between the lower bands $\Delta^{l}(\mathbf{k})=E^{l}_{-1}-E^{l}_{+1}$ in Fig. 3a and Fig. 3b, respectively. We also plot the dispersion along a path with high-symmetry points (Fig. 3c). From these we can confirm that we have an opposite chirality split along two different directions.
@@
The spin Nernst coefficient $\sigma_{\perp}$ exhibits a sinusoidal angular dependence and vanishes at $\theta=0,~{}\pi/2,~{}\pi$, where the thermal gradient is aligned with the crystal axes. At these angles, the transverse projection of the spin current cancels due to geometric symmetry. The extrema near $\theta=\pi/4$ and $43\pi/$ 4 arise when the thermal gradient is oriented diagonally, allowing both $\sigma_{xx}$ and $\sigma_{yy}$ to contribute significantly to the transverse response. Since the off-diagonal term $\sigma_{xy}$ is more than an order of magnitude smaller than the diagonal terms, the shape of $\sigma_{\perp}(\theta)$ is primarily governed by the anisotropy between $\sigma_{xx}$ and $\sigma_{yy}$, with $\sigma_{xy}$ introducing only a small asymmetry between the maximum and minimum.
@@ IV Conclusion
In this work, we have demonstrated that the $S=1$ Shastry-Sutherland model in the collinear Néel (I) phase exhibits altermagnetic behavior, characterized by magnon bands of opposite chiralities that are non-degenerate across the Brillouin zone. Through linear spin-wave analysis, we identified a unique pattern of chirality splitting, distinctively opposite along orthogonal directions, arising purely from the symmetry of the model, without requiring spin-orbit coupling or external magnetic fields.

### v2

@@ I Introduction
In this work, we demonstrate that the collinear Néel phase of the $S=1$ Shastry-Sutherland model possesses altermagnetic character. We establish this both from symmetry considerations and from a direct analysis of magnon dispersions. Our study reveals a unique form of chirality-splitting in the magnon bands: the splitting reverses sign along two different directions in momentum space. Furthermore, we compute the resulting spin Seebeck and spin Nernst coefficients using the Kubo formalism and show that these effects remain robust without the need for spin-orbit coupling or applied magnetic fields. Our findings enrich the growing landscape of altermagnetic phenomena and suggest new pathways for controlling magnonic transport in antiferromagnetic systems.
@@ II Model and Methods
where $\mathbf{S}_{i}$ denotes the spin operator on the $i$th site. As illustrated in Fig. 1a $_{ij}=J, JJ^{\prime}$, and $J^{\prime\prime}$ represent the intra-dimer, the inter-dimer and the inter-chain couplings, which are all assumed to be antiferromagnetic. The single-ion anisotropy is denoted as $D<0$.
@@
The Néel (I) phase (Fig. 1b) of the orthogonal-dimer structure belongs to the plane group $p4g$: the 4-fold axes are at the center and corner of the unit cell, and the 2-fold axes (inversion centers) are at the center of the edges of the cell. We verify the Néel (I) phase to be altermagnetic by checking the identification rules [19]:
@@
where the Nambu spinor $\Psi_{\mathbf{k}}=\left({b^{1}_{\mathbf{k}}},~{b^{2}_{\mathbf{k}}},~{b^{3}_{\mathbf{k}}},~{b^{4}_{\mathbf{k}}};~{b_{-\mathbf{k}}^{1}}^{\dagger},~{b_{-\mathbf{k}}^{2}}^{\dagger},~{b_{-\mathbf{k}}^{3}}^{\dagger},~{b_{-\mathbf{k}}^{4}}^{\dagger}\right)^{T}$, with $b^{l}_{\mathbf{k}}$$^{\dagger}$ (l=1,2,3,4)(l=1,~2,~3,~4) being the creation operator defined at the four sites in the unit cell. The Hamiltonian matrix in the reciprocal space takes the form of
@@
where $M(\mathbf{k})$ and $N(\mathbf{k})$ are $4\times 4$ Hermitian matrices:
@@
Figure 2: Magnon band dispersion of the generalized Shastry-Sutherland model near $\mathbf{k}=0$. Red and blue indicate the left and right-handed magnon modes, respectively. The values $^{\prime}/J=1.2, J′′/J$$^{\prime\prime}/J=1$, and D/J=0.6D/J=-0.6 are used for the parameters.
A distinct chirality, corresponding to the direction of precession of the spin around the axis of the collinear magnetic order, can be assigned to each of the magnon bands. Since the collinear order in the $z$-direction preserves the $z$-component of the total spin $S_{z}=\sum_{i}S_{i}^{z}$ with $i$ running over all sites, we use $\vartheta=\langle S^{z}\rangle/\hbar$ to specify the chirality of the bands [23]. In particular, we have $\vartheta=-1$ for the right-handed and $\vartheta=+1$ for left-handed magnon modes. In other words, magnon modes with chirality $\vartheta=\pm 1$ carry $\pm 1$ spin angular momentum along the $z$-direction. Since there are distinct upper and lower bands in the dispersion with opposite chiralities in each, we label the bands as $E^{u}_{\pm 1}(\mathbf{k})$ and $E^{l}_{\pm 1}(\mathbf{k})$, respectively, where the subscript denotes the chirality $\vartheta$ of the band (Fig. 2).
@@
where $\theta$ is the angle between the temperature gradient $\nabla T$ and the $\hat{\mathbf{x}}$ direction. Based on the Kubo formula [24, 25, 26, 27, 28, 29], we can calculate the magnon conductivity tensor $\sigma_{mn}=\sum_{\mu}\sigma_{mn}^{\mu}$ where $\sigma_{mn}^{\mu}$ is the contribution from the mode $\mu\in\{E^{u}_{\pm 1},~E^{l}_{\pm 1}\}$
@@
where $\bm{\sigma}$ is the original thermal conductivity tensor, and $\bm{\sigma}^{\prime}$ gives the components in the rotated frame. In this rotated frame, the diagonal component $\sigma^{\prime}_{x^{\prime}x^{\prime}}$ corresponds to the spin Seebeck response (along the gradient) [30], and the off-diagonal component $\sigma^{\prime}_{y^{\prime}x^{\prime}}$ corresponds to the spin Nernst response (transverse to the gradient) [31, 32, 33].
@@ III Results
Figure 3: (a) Energy difference between upper bands with right and left-handed chirality, (b) energy difference between lower bands with right and left-handed chirality, (c) chirality-split dispersion relation of magnon along the path $\Gamma\to M\to X\to\Gamma\to M\to X$ shown in the inset. We first trace the path $\Gamma\to M\to X\to\Gamma$ (blue arrows), and then $\Gamma\to M\to X$ (red arrows). Red and blue denote the left and right-handed magnon modes, respectively. Here we used $^{\prime}/J=1.2, J′′/J$$^{\prime\prime}/J=1$, and D/J=0.6D/J=-0.6. (d) Variation of spin Seebeck ($\sigma_{\parallel}$) and spin Nernst ($\sigma_{\perp}$) coefficients for different values of $\theta$. Here, in addition, we used $J/k_{B}=100\text{ K}$ and $T=25\text{ K}$.
The magnon dispersion is shown in Fig. 2. We have used the parameter values $^{\prime}/J=1.2, J′′/J$$^{\prime\prime}/J=1$, and D/J=0.6D/J=-0.6 unless otherwise stated. There are four bands: two with opposite chiralities with higher energy, $E^{u}_{\pm 1}$, and two with opposite chiralities with lower energy, $E^{l}_{\pm 1}$, with a strict energy gap between these two sectors. We will refer to these two pairs of bands as the lower and the upper bands, respectively.
We observe a $C_{2}$ symmetry in all of the magnon bands, as expected. To observe the chirality split, we calculated the energy difference between the upper bands $\Delta^{u}(\mathbf{k})=E^{u}_{-1}-E^{u}_{+1}$ and between the lower bands $\Delta^{l}(\mathbf{k})=E^{l}_{-1}-E^{l}_{+1}$ in Fig. 3a and Fig. 3b, respectively. We also plot the dispersion along a path with high-symmetry points (Fig. 3c). From these we can confirm that we have opposite chirality split along two different directions.
@@
The spin Nernst coefficient $\sigma_{\perp}$ exhibits a sinusoidal angular dependence and vanishes at $\theta=0,~\pi/2,~\pi$, where the thermal gradient is aligned with the crystal axes. At these angles, the transverse projection of the spin current cancels due to geometric symmetry. The extrema near $\theta=\pi/4$ and $3\pi/4$ arise when the thermal gradient is oriented diagonally, allowing both $\sigma_{xx}$ and $\sigma_{yy}$ to contribute significantly to the transverse response. Since the off-diagonal term $\sigma_{xy}$ is more than an order of magnitude smaller than the diagonal terms, the shape of $\sigma_{\perp}(\theta)$ is primarily governed by the anisotropy between $\sigma_{xx}$ and $\sigma_{yy}$, with $\sigma_{xy}$ introducing only a small asymmetry between the maximum and minimum.
@@ IV Conclusion
In this work, we have demonstrated that the $S=1$ Shastry-Sutherland model in the collinear Néel (I) phase exhibits altermagnetic behavior, characterized by magnon bands of opposite chiralities that are non-degenerate across the Brillouin zone. Through linear spin-wave analysis, we identified a unique pattern of chirality splitting, distinctively opposite along different directions, arising purely from the symmetry of the model, without requiring spin-orbit coupling or external magnetic fields.